The percentage ionic character of a compound AB is 20. Its dipole moment is 1.2D.The bond length in metre is

Grade:12

1 Answers

Shramana Kar
7 years ago
If the compound will be 100% ionic then it`s e=4.8*10^-10 esu & it`s dipole moment will be, u=4.8*10^-10 *d(d=the bond length in cm) esu.cm.So its ionic character,20=(1.2*100)/(4.8*10^8 *d) [ 1D=10^-18 esu.cm].hence the bond length ,d=1.5*10^-10 c.m=1.5*10^-12m.